# Detection of pairwise kSZ effect with DESI galaxy groups and Planck in Fourier space

### 07 Jan 2024 v1

@@ 3.2 Density-weighted pairwise kSZ power spectrum
The trivial average of the kSZ signal in eq. (3) vanishes due to its directionl dependence. Fortunately, under the influence of gravity, there is a tendency of two groups moving towards each other on quasi- to non-linear scales. In turn we can detect the kSZ effect by a pairwise estimator (Sugiyama et al., 2018a):
@@
here $\xi_{\rm kSZ}$ is the density-weighted pairwise kSZ correlation function, $\bm{s}_{ij}=\bm{s}_{i}-\bm{s}_{j}$ where $\bm{s}_{i}$ denotes the redshift space position of the $i$th group, $N$ represents the number of observed groups, and $V$ is the volume of the survey.
@@
Inserting eq. (3) into eqs. (4) and (5), $\xi_{\rm kSZ}$ and $P_{\rm kSZ}$ are proportional to the density-weighted pairwise velocity correlation function $\xi_{\rm pv}$ and power spectrum $P_{\rm pv}$ respectively:
@@
Compared to the traditional pair-weighted statistics, e.g. $\xi_{\rm pv}^{\rm pair-weight}=\xi_{\rm pv}/(1+\xi_{\delta})$, the density-weighted statistics encode similar cosmological information, while their modelling is simpler by avoiding the modelling of an additional density correlation function $\xi_{\delta}$ in the denominator. A detailed comparison between the density- and pair-weighted statistics can be found in the appendix B of Okumura et al. (2014).
@@
In this work, we use the following non-linear galaxy power spectrum to derive the kSZ power spectrum:
@@
In order to tackle the non-linear correlation between density and velocity fields, we include the window function $\tilde{W}(k)\equiv P_{\delta\theta}/fP_{\delta\delta}$ in eq. (11(Zhang et al., 2013). Motivated by the perturbation theory and verified with N-body simulations, $\tilde{W}(k)$ can be described by the following fitting function (Zheng et al., 2013)
@@
By inserting eq. (11) into eqs. (7) and (10), we derive the non-linear density-weighted pairwise kSZ power spectrum as
@@
At linear scales, the above non-linear model can be simplified to a linear one
@@
It is easy to figure out the degeneracy between $b_{1}$ and $\bar{\tau}$, namely that an accurate $\bar{\tau}$ detection requires an accurate $b_{1}$ constraints. In the next section, we further notice that the photo-z uncertainty heavily damps the kSZ power spectrum, and this damping is also degenerated with $b_{1}$ and $\bar{\tau}$. This is a unique feature for the photo-z data. In appendix A, we tackle these issues by applying a joint analysis of kSZ and RSD analysis on the same dataset. Tests there show that our $\bar{\tau}$ measurements are minimally biased by the above mentioned degeneacies.
@@ 3.3 Density-weighted pairwise kSZ power spectrum with photometric redshits
Figure 3: Theoretical predictions of the unmasked pairwise kSZ multipoles of the NGC sub-sample of the baseline catalog. Solid lines represent multipoles of the non-linear density-weighted pairwise kSZ power spectrum with photo-z errors (eq. (17)) and dashed lines are those without photo-z errors (eq. (14)). The red and blue lines are for dipoles ($\ell=1$) and octopoles ($\ell=3$) respectively. The adopted redshift is the median redshift $\tilde{z}=0.64$ of the NGC sub-sample, and $\bar{\tau}=3.96\times 10^{-5}$ is the best fitted $\bar{\tau}$ value of the baseline sample.
The photometric data we analyze have non-negligible redshift errors. These photo-z errors randomly distort the LOS positions of galaxy groups and dilute their LOS clustering, which is an analogy of the FoG effect. As discussed in detail in appendix A, we model the distribution function of the pairwise photo-z errors as an exponential function with zero mean and a dispersion of $\sigma^{\rm pair}_{\rm pho-z}=\sqrt{2}\sigma_{\rm pho-z}$, with $\sigma_{\rm pho-z}$ being the dispersion of the 1-D photo-z error distribution function, then the galaxy power spectrum of the photo-z data becomes (Davis & Peebles, 1983; Ballinger et al., 1996)
@@
In appendix A, we measure the galaxy density power spectrum multipoles of group catalogs and fit the measurements with the non-linear model in eqs. (11) and (16). With $f$ being fixed to the fiducial cosmological value, this fitting determines the posterior distribution of $b_{1}$, $b_{2}$ and $\sigma_{z}$. We then substitute the best fitted values of these 3 parameters to eq. (14) and fit the kSZ measurement to obtain the constraint on $\bar{\tau}$. In this way, (1) we model the galaxy density and kSZ power spectra in a self-consistent manner; (2) we self-consistently break the $b_{1}-\bar{\tau}$ degeneracy and avoid possible systematic biases coming from a manually selected $b_{2}-M$ relation and $\sigma_{\rm pho-z}$.; (3) as $\bar{\tau}$ is the only parameter in the kSZ measurement fitting, the S/N of the fitted $\bar{\tau}$ directly represents the S/N of the kSZ detection, which is one of main targets of this work. For a global fitting of the galaxy density and kSZ power spectra simultaneously, which requires calculations of the covariance matrix between two power spectra, we defer it to the future work.
@@
Since $P_{\rm kSZ}(k,\mu)$ is an odd function on $\mu$, $P^{\ell}_{\rm kSZ}$ is non-zero only for odd $\ell$’s, in contrast to the galaxy power spectrum $P_{\rm s}(k,\mu)$. As shown by solid lines in figure 3, the absolute values of $P^{\ell=3}_{\rm kSZ}(k,\mu)$ are smaller than those of $P^{\ell=1}_{\rm kSZ}(k,\mu)$, and in principle the measurement errors of the former will be larger. Considering the limited S/N of our measurement, we only focus on the dipole $P^{\ell=1}_{\rm kSZ}$ in this work.
By dashed lines, we also plot the non-linear kSZ power spectrum multipoles without photo-z errors evaluated by eq. (14) in figure 3. These will be signals likely detected by future spectroscopic galaxy surveys, e.g., DESI (DESI Collaboration et al., 2016) or PFS (Takada et al., 2014). Without the smearing effect of photo-z errors, kSZ multipoles exhibit more measurable features at non-linear scales, and will guarantee a kSZ detection of higher S/N (Sugiyama et al., 2018a).
@@ 4.1 Aperture photometry filter
In order to minimize the contamination from the primary CMB, an AP filter is applied to the Planck CMB map (e.g. Sugiyama et al. (2018a); Calafut et al. (2021); Chen et al. (2022)). The AP filter is a 2-D compensated top-hat filter:
@@
In this work, we follow (Chen et al., 2022) and apply the AP filter in the spherical harmonic space, namely
@@ 4.2 Random catalog
We rely on the random catalog to calculate the galaxy group density field in eq. (25). We generate a corresponding random catalog for each group sample we construct, by randomly downsizing the random catalog provided along with the DESI Legacy Imaging Survey DR9 data [^7] https://portal.nersc.gov/cfs/cosmo/data/legacysurvey/dr9/randoms/. The size of each random catalog is 20 times of the size of the corresponding group sample. Since the group catalog of Yang et al. (2021) is constructed within the continuous region of the DESI sky coverage, we manually discard the discontinuous regions in the random catalog in the first step. The pixelized sky distribution of the random catalog is shown on the right panel of figure 1.
@@
In section 4.1, the AP filtered CMB temperature $\Delta T_{\rm kSZ}^{\rm AP}$ of each group is estimated in the spherical harmonic space. To suppress the systematic errors induced by the redshift-dependent foregrounds such as the cosmic infrared background (CIB) and the residual tSZ signal, we subtract from $\Delta T_{\rm kSZ}^{\rm AP}$ its average with a Gaussian weight (Hand et al., 2012):
@@
where $G(z_{i},z_{j},\sigma_{z})={\rm exp}(-z^{2}_{ij}/2\sigma^{2}_{z})$ with $\sigma_{z}=0.01$, and $z_{ij}\equiv z_{i}-z_{j}$. The 3-D kSZ temperature fluctuation field then reads
@@
where $N_{\rm group}$ is the size of the group sample and $\delta_{\rm D}$ is the 3-D Dirac delta function.
With the help of the random catalog constructed in section 4.2, we define the 3-D galaxy group density field as
@@ 4.4 Estimator of the pairwise kSZ power spectrum multipoles
We adopt the estimator of the density-weighted pairwise kSZ power spectrum multipoles developed by (Sugiyama et al., 2018a), which is an analogy to the estimator of the galaxy power spectrum multipoles (Feldman et al., 1994; Yamamoto et al., 2005),
@@
where the number density $\bar{n}$ is directly measured from the random catalogue. We do not use the FKP weight in this estimator since it is not derived for an optimal kSZ detection.
The shot noise term disappears in eq. (26), as it is cancelled out by the subtraction of the two terms in the square bracket. So do the detector noise and the primary CMB anisotropies contaminating the $\delta T(\bm{s})$ measurement. The pairwise estimator guarantees that they will not bias the kSZ power spectrum, although their contributions to the covariance matrix remain.
@@
eq. (26) can be simplified as (Yamamoto et al., 2005; Sugiyama et al., 2018a)
@@
The evaluation of the Legendre polynomial involved quantity ($\delta n_{\ell}$ here) is time consuming. It is proposed in Bianchi et al. (2015); Scoccimarro (2015) that by decomposing $\hat{\bm{k}}\cdot\hat{\bm{s}}$ into its Cartesian components, eq. (30) can be expressed as a sum over the Fourier transforms of $\delta n(\bm{s})$ weighted by products of Cartesian vectors. Thus the application of Fast Fourier Transform (FFT) technique888numpy.fft.fftn can speed up the calculation in several orders of magnitude as compared to a naive summation implementation of eq. (30). The number of FFTs required in each $\hat{P}_{\ell}$ calculation is $(\ell+1)(\ell+2)/2$ in this scenario.
@@
In this case, implication of only $12\ell+$ 1 times of FFTs are needed to evaluate each $\delta n_{\ell}$. Although when $\ell=1$, both methods require 3 FFTs, the improvement will become evident as $\ell$ increases. Therefore we adopt the second method to calculate the pairwise kSZ power spectrum dipole, and the calculation of eq. (33) heavily refers to the software toolkit nbodykit (Hand et al., 2018).
Finally, we select appropriate $k$ bins, $k_{i}\in[0.05,0.155]\ (h/\rm Mpc)$ with interval $\Delta k=0.01h/\rm Mpc$, and average the $|\bm{k}|$ valves within each bin to obtain an effective $k$ value. This tackles the binning discreteness effect, which can be prominent at linear scales. Being dominated by noises and photo-z damping, larger $k$ modes can not increase the measurement S/N. Furthermore, we average the measured $\hat{P}^{\ell}_{\rm kSZ}$ within each $k$ bin from eq. (29) to compute our estimated $\hat{P}_{\rm kSZ}^{\ell}(k)$:
@@ 4.5 Measurement details
where $A_{\rm NGC}$ and $A_{\rm SGC}$ are effective coverage of the contiguous areas of NGC and SGC, namely $A_{\rm NGC}=9622\ \rm deg^{2}$ and $A_{\rm SGC}=8601\ \rm deg^{2}$. Eq. (36) is used analogically to estimate other parameters, such as the median redshift, median mass of the whole group sample, etc.
The FFTs are performed on regular grids with a fiducial size of $512^{3}$. We establish the Cartesian coordinates $\bm{s}=(x,y,z)$ with the $z$ axis pointing to the north pole. Then we locate the group samples into a cuboid of dimensions $(L_{x},L_{y},L_{z})$, with $(L_{x},L_{y},L_{z})[h^{-1}\rm Mpc]=(2400,4000,2600)$ for NGC, and $(2600,4000,3500)$ for SGC. The averaged grid size is $\sim 6h^{-1}\rm Mpc$ for both NGC and SGC. The triangular-shaped cloud (TSC) assignment function is used to distribute both the groups and associated $\delta T$’s onto the grids. This gridding process induces numerical artifacts like the aliasing effect and the window function effect (e.g., Jing (2005)). In Fourier space, we follow the interlacing technique proposed by Hockney & Eastwood (1981); Sefusatti et al. (2016) to alleviate the aliasing effect, and correct the window function (e.g., eq. (18) of Jing (2005)) for each $\bm{k}$ mode on the grids. As illustrated by figure 2 of Hand et al. (2018), these corrections guarantee to reduce the corresponding artifacts down to a sub-percent level even at scales near the Nyquist frequency in our case and will not bias the results significantly. We also try the measurement on a $1024^{3}$ grid and little difference can be observed within the $k$ range we use. More details about the FFT related procedure can be found in appendix B2 of Sugiyama et al. (2018a).

### 01 Mar 2024 v2

@@ 3.2 Density-weighted pairwise kSZ power spectrum
The trivial average of the kSZ signal in Equation (3) vanishes due to its directionl dependence. Fortunately, under the influence of gravity, there is a tendency of two groups moving toward each other on quasi to nonlinear scales. In turn, we can detect the kSZ effect by a pairwise estimator (Sugiyama et al., 2018a):
@@
where $\xi_{\rm kSZ}$ is the density-weighted pairwise kSZ correlation function, $\bm{s}_{ij}=\bm{s}_{i}-\bm{s}_{j}$ where $\bm{s}_{i}$ denotes the redshift space position of the $i$th group, $N$ represents the number of observed groups, and $V$ is the volume of the survey.
@@
Inserting Equation (3) into Equations (4) and (5), $\xi_{\rm kSZ}$ and $P_{\rm kSZ}$ are proportional to the density-weighted pairwise velocity correlation function $\xi_{\rm pv}$ and power spectrum $P_{\rm pv}$ respectively:
@@
Compared to the traditional pair-weighted statistics, e.g., $\xi_{\rm pv}^{\rm pair-weight}=\xi_{\rm pv}/(1+\xi_{\delta})$, the density-weighted statistics encode similar cosmological information, while their modeling is simpler by avoiding the modeling of an additional density correlation function $\xi_{\delta}$ in the denominator. A detailed comparison between the density- and pair-weighted statistics can be found in Appendix B of Okumura et al. (2014).
@@
In this work, we use the following nonlinear galaxy power spectrum to derive the kSZ power spectrum:
@@
In order to tackle the nonlinear correlation between density and velocity fields, we include the window function $\tilde{W}(k)\equiv P_{\delta\theta}/fP_{\delta\delta}$ in Equation (11(Zhang et al., 2013). Motivated by the perturbation theory and verified with N-body simulations, $\tilde{W}(k)$ can be described by the following fitting function (Zheng et al., 2013)
@@
By inserting Equation (11) into Equations (7) and (10), we derive the nonlinear density-weighted pairwise kSZ power spectrum as
@@
At linear scales, the above nonlinear model can be simplified to a linear one
@@
It is easy to figure out the degeneracy between $b_{1}$ and $\bar{\tau}$, namely that an accurate $\bar{\tau}$ detection requires an accurate $b_{1}$ constraint. In the next section, we further notice that the photo-z uncertainty heavily damps the kSZ power spectrum, and this damping is also degenerated with $b_{1}$ and $\bar{\tau}$. This is a unique feature for the photo-z data. In Appendix A, we tackle these issues by applying a joint analysis of kSZ and RSD analysis on the same data set. The tests there show that our $\bar{\tau}$ measurements are minimally biased by the abovementioned degeneacies.
@@ 3.3 Density-weighted pairwise kSZ power spectrum with photometric redshits
Figure 3: Theoretical predictions of the unmasked pairwise kSZ multipoles of the NGC sub-sample of the baseline catalog. Solid lines represent multipoles of the nonlinear density-weighted pairwise kSZ power spectrum with photo-z errors (Equation (17)) and dashed lines are those without photo-z errors (Equation (14)). The red and blue lines are for dipoles ($\ell=1$) and octopoles ($\ell=3$) respectively. The adopted redshift is the median redshift $\tilde{z}=0.64$ of the NGC sub-sample, and $\bar{\tau}=3.96\times 10^{-5}$ is the best-fitted $\bar{\tau}$ value of the baseline sample.
The photometric data we analyze have nonnegligible redshift errors. These photo-z errors randomly distort the LOS positions of galaxy groups and dilute their LOS clustering, which is an analogy of the FoG effect. As discussed in detail in Appendix A, we model the distribution function of the pairwise photo-z errors as an exponential function with zero mean and a dispersion of $\sigma^{\rm pair}_{\rm pho-z}=\sqrt{2}\sigma_{\rm pho-z}$, with $\sigma_{\rm pho-z}$ being the dispersion of the 1-D photo-z error distribution function; then, the galaxy power spectrum of the photo-z data becomes (Davis & Peebles, 1983; Ballinger et al., 1996)
@@
In Appendix A, we measure the galaxy density power spectrum multipoles of group catalogs and fit the measurements with the nonlinear model in Equations (11) and (16). With $f$ being fixed to the fiducial cosmological value, this fitting determines the posterior distribution of $b_{1}$, $b_{2}$ and $\sigma_{z}$. We then substitute the best-fitted values of these 3 parameters to Equation (14) and fit the kSZ measurement to obtain the constraint on $\bar{\tau}$. In this way, (1) we model the galaxy density and kSZ power spectra in a self-consistent manner; (2) we self-consistently break the $b_{1}-\bar{\tau}$ degeneracy and avoid possible systematic biases coming from a manually selected $b_{1}-M$ relation and $\sigma_{\rm pho-z}$.; (3) as $\bar{\tau}$ is the only parameter in the kSZ measurement fitting, the S/N of the fitted $\bar{\tau}$ directly represents the S/N of the kSZ detection, which is one of main targets of this work. For a global fitting of the galaxy density and kSZ power spectra simultaneously, which requires calculations of the covariance matrix between two power spectra, we defer it to the future work.
@@
Since $P_{\rm kSZ}(k,\mu)$ is an odd function on $\mu$, $P^{\ell}_{\rm kSZ}$ is nonzero only for odd $\ell$s, in contrast to the galaxy power spectrum $P_{\rm s}(k,\mu)$. As shown by solid lines in figure 3, the absolute values of $P^{\ell=3}_{\rm kSZ}(k,\mu)$ are smaller than those of $P^{\ell=1}_{\rm kSZ}(k,\mu)$, and in principle, the measurement errors of the former will be larger. Considering the limited S/N of our measurement, we only focus on the dipole $P^{\ell=1}_{\rm kSZ}$ in this work.
By dashed lines, we also plot the nonlinear kSZ power spectrum multipoles without photo-z errors evaluated by Equation (14) in figure 3. These will be signals likely detected by future spectroscopic galaxy surveys, e.g., DESI (DESI Collaboration et al., 2016) or Subaru Prime Focus Spectrograph (PFS) (Takada et al., 2014). Without the smearing effect of photo-z errors, kSZ multipoles exhibit more measurable features at nonlinear scales, and will guarantee a kSZ detection of higher S/N (Sugiyama et al., 2018a).
@@ 4.1 Aperture photometry filter
In order to minimize the contamination from the primary CMB, an AP filter is applied to the Planck CMB map (e.g. Sugiyama et al. (2018a); Calafut et al. (2021); Chen et al. (2022)). The AP filter is a 2D compensated top-hat filter:
@@
In this work, we follow Chen et al. (2022) and apply the AP filter in the spherical harmonic space, namely
@@ 4.2 Random catalog
We rely on the random catalog to calculate the galaxy group density field in Equation (25). We generate a corresponding random catalog for each group sample we construct, by randomly downsizing the random catalog provided along with the DESI Legacy Imaging Survey DR9 data [^7] https://portal.nersc.gov/cfs/cosmo/data/legacysurvey/dr9/randoms/. The size of each random catalog is 20 times of the size of the corresponding group sample. Since the group catalog of Yang et al. (2021) is constructed within the continuous region of the DESI sky coverage, we manually discard the discontinuous regions in the random catalog in the first step. The pixelized sky distribution of the random catalog is shown on the right panel of figure 1.
@@
In Section 4.1, the AP filtered CMB temperature $\Delta T_{\rm kSZ}^{\rm AP}$ of each group is estimated in the spherical harmonic space. To suppress the systematic errors induced by the redshift-dependent foregrounds such as the cosmic infrared background and the residual tSZ signal, we subtract from $\Delta T_{\rm kSZ}^{\rm AP}$ its average with a Gaussian weight (Hand et al., 2012):
@@
where $G(z_{i},z_{j},\sigma_{z})={\rm exp}(-z^{2}_{ij}/2\sigma^{2}_{z})$ with $\sigma_{z}=0.01$, and $z_{ij}\equiv z_{i}-z_{j}$. The 3D kSZ temperature fluctuation field then reads
@@
where $N_{\rm group}$ is the size of the group sample and $\delta_{\rm D}$ is the 3D Dirac delta function.
With the help of the random catalog constructed in Section 4.2, we define the 3D galaxy group density field as
@@ 4.4 Estimator of the pairwise kSZ power spectrum multipoles
We adopt the estimator of the density-weighted pairwise kSZ power spectrum multipoles developed by Sugiyama et al. (2018a), which is an analogy to the estimator of the galaxy power spectrum multipoles (Feldman et al., 1994; Yamamoto et al., 2005),
@@
where the number density $\bar{n}$ is directly measured from the random catalog. We do not use the FKP weight in this estimator since it is not derived for an optimal kSZ detection.
The shot noise term disappears in Equation (26), as it is canceled out by the subtraction of the two terms in the square bracket. This is also true for the detector noise and the primary CMB anisotropies contaminating the $\delta T(\bm{s})$ measurement. The pairwise estimator guarantees that they will not bias the kSZ power spectrum, although their contributions to the covariance matrix remain.
@@
Equation (26) can be simplified as (Yamamoto et al., 2005; Sugiyama et al., 2018a)
@@
The evaluation of the Legendre polynomial involved quantity ($\delta n_{\ell}$ here) is time consuming. It is proposed in Bianchi et al. (2015); Scoccimarro (2015) that, by decomposing $\hat{\bm{k}}\cdot\hat{\bm{s}}$ into its Cartesian components, Equation (30) can be expressed as a sum over the Fourier transforms of $\delta n(\bm{s})$ weighted by products of Cartesian vectors. Thus, the application of FFT technique888numpy.fft.fftn can speed up the calculation in several orders of magnitude as compared to a naive summation implementation of Equation (30). The number of FFTs required in each $\hat{P}_{\ell}$ calculation is $(\ell+1)(\ell+2)/2$ in this scenario.
@@
In this case, the implication of only $12\ell+$ 1 times of FFTs is needed to evaluate each $\delta n_{\ell}$. Although when $\ell=1$, both methods require 3 FFTs; the improvement will become evident as $\ell$ increases. Therefore, we adopt the second method to calculate the pairwise kSZ power spectrum dipole, and the calculation of Equation (33) heavily refers to the software toolkit nbodykit (Hand et al., 2018).
Finally, we select appropriate $k$ bins, $k_{i}\in[0.05,0.155]\ (h/\rm Mpc)$ with interval $\Delta k=0.01h/\rm Mpc$, and average the $|\bm{k}|$ valves within each bin to obtain an effective $k$ value. This tackles the binning discreteness effect, which can be prominent at linear scales. Being dominated by noises and photo-z damping, larger $k$ modes cannot increase the measurement S/N. Furthermore, we average the measured $\hat{P}^{\ell}_{\rm kSZ}$ within each $k$ bin from Equation (29) to compute our estimated $\hat{P}_{\rm kSZ}^{\ell}(k)$:
@@ 4.5 Measurement details
where $A_{\rm NGC}$ and $A_{\rm SGC}$ are effective coverage of the contiguous areas of NGC and SGC, namely $A_{\rm NGC}=9622\ \rm deg^{2}$ and $A_{\rm SGC}=8601\ \rm deg^{2}$. Equation (36) is used analogically to estimate other parameters, such as the median redshift, median mass of the whole group sample, etc.
The FFTs are performed on regular grids with a fiducial size of $512^{3}$. We establish the Cartesian coordinates $\bm{s}=(x,y,z)$ with the $z$-axis pointing to the north pole. Then, we locate the group samples into a cuboid of dimensions $(L_{x},L_{y},L_{z})$, with $(L_{x},L_{y},L_{z})[h^{-1}\rm Mpc]=(2400,4000,2600)$ for NGC, and $(2600,4000,3500)$ for SGC. The averaged grid size is $\sim 6h^{-1}\rm Mpc$ for both NGC and SGC. The triangular-shaped cloud (TSC) assignment function is used to distribute both the groups and associated $\delta T$s onto the grids. This gridding process induces numerical artifacts like the aliasing effect and the window function effect (e.g., Jing (2005)). In Fourier space, we follow the interlacing technique proposed by Hockney & Eastwood (1981); Sefusatti et al. (2016) to alleviate the aliasing effect, and correct the window function (e.g., Equation (18) of Jing (2005)) for each $\bm{k}$ mode on the grids. As illustrated by figure 2 of Hand et al. (2018), these corrections guarantee to reduce the corresponding artifacts down to a subpercent level even at scales near the Nyquist frequency in our case and will not bias the results significantly. We also try the measurement on a $1024^{3}$ grid, and little difference can be observed within the $k$ range we use. More details about the FFT related procedure can be found in Appendix B2 of Sugiyama et al. (2018a).